Write an equation in point-slope form for the line through the given point with the given slope: (-10, 9) and m = -2
Answer by jim_thompson5910(35256)

y - y1 = m(x - x1)

y - 9 = -2(x - (-10))

y - 9 = -2(x + 10)

So the equation in point slope form is y - 9 = -2(x + 10)